OpenStudy (anonymous):

Solve the following trig equation for 0°≤θ<360° -9cosθ+5=6cosθ+1

OpenStudy (dls):

Rearrange the equation to get : 15 cos theta = 4 and solve for theta

OpenStudy (anonymous):

=>16 cos theta=4 => cos theta= 1/4 =>theta= cos^-1(1/4) =>theta=75.52

OpenStudy (dls):

9+6=15

OpenStudy (dls):

my dear human calculator :)

OpenStudy (dumbcow):

you will need to use your calculator to do inverse cos \[\theta = \cos^{-1} \frac{4}{15}\] Note you will have 2 solutions in 1st and 4th quadrants because cos is positive

OpenStudy (dumbcow):



OpenStudy (anonymous):

like i don't understand the whole proccess :$

OpenStudy (dumbcow):

which part?

OpenStudy (anonymous):

how is there 16 and 15 in the first 2 answers?

OpenStudy (dumbcow):

from using algebra to combine terms and isolate the cos ignore the 16, it was a typo or simple mistake in adding

OpenStudy (anonymous):

huh okay, then how about this equation: Solve the following trig equation for 0°≤θ<360° -4cos2θ=-3

OpenStudy (anonymous):

i don't really understand the general idea so maybe another example may be helpful @dumbcow :4